# From Cursed to Competitive: Closing the ZO–FO Gap via Input-to-State Stability

### 28 Apr 2026 v1

@@ 4.2 Binary classification with a neural network
Figure 2 shows the training loss for all six methods (GD, HB, NAG and their ZO counterparts). We set $h=1\times 10^{-3}$ for GD and ZO-GD, $h_{1}=1\times 10^{-3}$, $h_{2}=5\times 10^{-4}$ for HB/NAG and their ZO counterparts with smoothing parameter $\mu=10^{-7}$ and number of sampled directions $t=7$. The ZO trajectories closely track their FO counterparts, and all methods reach $99.55\%$ training accuracy and $99.40\%$ test accuracy. It takes approximately 4 times longer to run the ZO methods compared to their FO counterpart. Parameter sensitivity experiments confirming the same phenomena as in the quadratic case are provided in Appendix E.2.
@@
To directly validate the absence of extra dimension dependence, we vary $H\in\{16,32,64,128,256\}$ ($n$ from ${\approx}12{,}700$ to ${\approx}201{,}000$) and run GD and ZO-GD with identical parameters for 1000 iterations. We set $h=5\times 10^{-4},$ $t=10,$ and $\mu=10^{-8}.$ Figure 3 (left) shows that ZO-GD tracks GD at the same rate across all dimensions. Figure 3 (right) confirms the per-iteration time ratio remains approximately constant. Table 1 reports iteration counts to reach within $20\%$ of the FO final loss: the ratio ZO/FO stays close to $1$, confirming no extra dimension-dependent factor in the iteration complexity.
@@
Table 1: Dimension scaling: iterations to reach within $20\%$ of FO final loss reduction.
@@ 5 Conclusion and future directions
In this work, we revisited the commonly held belief that ZO methods inherently suffer from additional dimension dependence in their iteration complexity compared to their FO counterparts. By adopting a dynamical systems perspective, we showed that, under suitable conditions, the averaged dynamics of ZO algorithms can be interpreted as perturbed versions of their FO counterparts with bounded, controllable perturbations. Leveraging ISS arguments, we proved that ZO methods inherit the same convergence rates as FO methods in expectation, converging to a neighbourhood of the FO fixed point whose size can be made arbitrarily small by tuning the algorithm parameters. These findings suggest that the practical efficiency of ZO methods may be closer to that of FO methods than previously understood, particularly when gradient information is unavailable or expensive.
@@ A.1 Comparison with prior work
Table 2 summarises the convergence guarantees of representative ZO methods and highlights how our framework differs from existing analyses. In all prior works, the iteration complexity carries an explicit factor depending on the problem dimension $n$ (or an effective dimension $n_{\mathrm{eff}}$), and convergence is to the exact optimiser. By contrast, our ISS-based framework eliminates the dimension factor from the iteration complexity entirely; the dimension dependence instead appears in the radius of the convergence neighbourhood, which can be made arbitrarily small by tuning the algorithm parameters $\mu$, $h$, and $t$. Moreover, our framework applies uniformly to ZO-GD, ZO-HB, and ZO-NAG, whereas each prior result requires a separate, algorithm-specific analysis.
@@
Reference Assumptions on $f$ Iteration Dim. dependence Convergence target
[28] convex / s.c. $O(n)$ factor Minimiser
[7] convex $O(n)$ factor Minimiser
[16] nonconvex $O(n)$ factor $\|\nabla f\|^{2}\leq\varepsilon$
[39] convex / s.c; Hessian regularity $O(n_{\mathrm{eff}})$, $n_{\mathrm{eff}}\leq n$ Minimiser
[35] convex; sparse gradients $O(s\log n)$, $s$-sparse Minimiser
This work s.c. (Thms. 3–5); nonconvex + $L_{2}$ reg. (Thm. 6) No explicit dependence Neighbourhood of FO fixed point (can be made arbitrarily small)
@@ Appendix B Additional details and proofs of Section 2
It remains to choose $\bar{r}>0$ and $\bar{q}>0$ so that $\{z_{k}\}_{k\in\mathbb{N}}\subset\mathcal{B}_{R}(z^{e})$. For $z\in\mathbb{R}^{n}$ such that $V(z)\leq\ell=\frac{R^{2}}{c_{2}}$, (6) implies that $\|z-z^{e}\|^{2}\leq R^{2}$, i.e., $x\in\mathcal{B}_{R}(z^{e})$, and we can conclude $\{z\in\mathbb{R}^{n}|\ V(z)\leq\frac{R^{2}}{c_{2}}\}\subset\mathcal{B}_{R}(z^{e})$. From (41), the set $\{z\in\mathbb{R}^{n}|\ V(z)\leq\frac{R^{2}}{c_{2}}\}$ is forward-invariant for the perturbed system whenever $\lambda\,\ell+L_{V}\bar{q}\leq\ell$, i.e. whenever $\bar{q}\leq c\,\ell/L_{V}$. Setting
$\bar{r}:=\frac{\ell}{c_{1}},\qquad\bar{q}:=\frac{c\,\ell}{L_{V}}$
@@
At first glance, $kd$ grows without bound. But recall that $v_{k}$ eventually enters and remains in $0,\sigma(d)]$. Let $K^{*}$ be the first time $v_{k}\leq\sigma(d)$; from that point on, $v_{k}\leq\sigma(d)$ for all subsequent times. The strict decrease above $\sigma(d)$ gives a minimum decrease per step: for $v_{k}\in[\sigma(d),\max(R,\sigma(d))]$, we have
@@
On the compact set $\sigma(d),\max(R,\sigma(d))]$, the continuous function $(\mathrm{Id}-\varphi)(s)-d$ achieves its minimum, which is strictly positive (being zero only at $s=\sigma(d)$, which is the left endpoint). Hence the entry time satisfies
$K^{*}\leq\left\lceil\frac{R-\sigma(d)}{\displaystyle\min_{s\in[\sigma(d),R]}\bigl((\mathrm{Id}-\varphi)(s)-d\bigr)}\right\rceil$
when $R>\sigma(d)$ (and $K^{*}=0$ otherwise). In particular, $K^{*}$ depends only on $R$ and $d$ (not on $v_{0}$ beyond the constraint $v_{0}\leq R$).
Now we bound the excess for all $k$:
  • For $k\leq K^{*}$: using $v_{k}\leq\max(R,\sigma(d))$ from (42) and $\hat{\beta}(v_{0},k)\geq 0$, we get $e_{k}\leq\max(R,\sigma(d))$. But more precisely, $e_{k}=v_{k}-u_{k}\leq K^{*}\cdot d$ by the iterated bound above.
  • For $k>K^{*}$: $v_{k}\leq\sigma(d)$ and $\hat{\beta}(v_{0},k)\geq 0$, so $e_{k}\leq\sigma(d)$.
@@
$e_{k}\leq\max\bigl(K^{*}\cdot d,\;\sigma(d)\bigr)=:\eta_{R}(d).$
Since $K^{*}$ depends only on $R$ and $d$, so does $\eta_{R}(d)$. Moreover, $\eta_{R}(0)=0$ because $K^{*}=0$ and $\sigma(0)=0$ when $d=0$. The function $\eta_{R}$ is non-negative, finite, nondecreasing in $d$, and satisfies $\eta_{R}(0)=0$. By regularisation, there exists $\hat{\gamma}_{R}\in\mathcal{K}$ with $\eta_{R}(d)\leq\hat{\gamma}_{R}(d)$ for all $d\geq 0$.
@@ Appendix C Proof of lemmas and theorems of Section 3
Using [28, Lemma 3], we know that $\|\nabla f(\bar{z}_{k})-\nabla f_{\mu}(\bar{z}_{k})\|\leq\tfrac{\mu}{2}L_{1}(n+3)^{3/2}.$ Thus, the first term on the right-hand side (46) is bounded and can be manipulated through the choice of $h$ and $\mu.$ For the second term on the right-hand side of (46), we have
@@
$E[\|e_{k+1}\|]\leq(1+2Lh)E[\|e_{k}\|]+h\frac{\sigma}{\sqrt{t}},$ (49)
@@
$\begin{split}E[\|l_{k}\|^{2}]&=E_{U_{k}}[\|\nabla f_{\mu}(\tilde{z}_{k})-\nabla f_{\mu}(\bar{z}_{k})+\nabla f_{\mu}(\bar{z}_{k})-E_{U_{k}}[\nabla f_{\mu}(\tilde{z}_{k})]\|^{2}]\\ &=E[\|\nabla f_{\mu}(\tilde{z}_{k})-\nabla f_{\mu}(\bar{z}_{k})\|^{2}+\|\nabla f_{\mu}(\bar{z}_{k})-E_{U_{k}}[\nabla f_{\mu}(\tilde{z}_{k})]\|^{2}]\\ &\qquad-2E[\langle\nabla f_{\mu}(\tilde{z}_{k})-\nabla f_{\mu}(\bar{z}_{k}),E_{U_{k}}[\nabla f_{\mu}(\tilde{z}_{k})]\nabla f_{\mu}(\bar{z}_{k})\rangle]\\ &\leq E[\|\nabla f_{\mu}(\tilde{z}_{k})-\nabla f_{\mu}(\bar{z}_{k})\|^{2}]\\ &\leq L_{1}^{2}E[\|e_{k}\|^{2}].\end{split}$ (53)
@@
For $\beta$-strongly convex $f$ with $L_{1}$-Lipschitz gradients and $h\in\bigl(0,\frac{2}{\beta+L_{1}}\bigr)$, from [27, Thm. 2.1.15] we have
@@
Using [28, Lem. 3], we know that $\|\nabla f_{\mu}(x)-\nabla f(x)\|\leq\frac{\mu}{2}\,L_{1}\,(n+3)^{3/2}$ uniformly in $x$. Thus, the first term in the right-hand side (56) is bounded and can be ensured to be arbitrarily small by selecting $h_{1}$ and $\mu$ appropriately. For the second term in the right-hand side of (56), by $L_{1}$-Lipschitz continuity of $\nabla f$ and Jensen’s inequality, we have
@@
Since we assume that the variance of the random oracle is bounded by $E[\|v_{k}\|^{2}]\leq\sigma^{2}/t$, for $l_{k}$, similar to (53), it holds that $E[\|l_{k}\|^{2}]\leq L_{1}^{2}\,a_{k}$ and thus $E[\|v_{k}+l_{k}\|^{2}]\leq\frac{2\sigma^{2}}{t}+8L_{1}^{2}\,a_{k}$.
@@
$\sup_{k}\,E[\|q_{k}\|]$ $\leq\frac{h_{1}\mu\,L_{1}(n+3)^{3/2}}{2}+h_{1}L_{1}\cdot\frac{h_{1}\sigma}{\sqrt{t}}\,\sqrt{\frac{2}{1-\rho_{\mathrm{HB}}}}$
@@
Using [28, Lem. 3], we know that $\|\nabla f_{\mu}(x)-\nabla f(x)\|\leq\frac{\mu}{2}L_{1}(n+3)^{3/2}$ uniformly in $x$. Thus, the first term in the right-hand side (73) is bounded and can be controlled by the choice of $h_{1}$ and $\mu$. For the second term in the right-hand side of (73), by $L_{1}$-Lipschitz continuity of $\nabla f$ and Jensen’s inequality, we have
@@
Choosing $h_{1}\in(0,\frac{\beta}{8L_{1}^{2}})$ ensures $\psi\in(0,1)$, and then any $h_{2}\in(0,\frac{1}{4}(\psi^{-1/2}-1))$ guarantees (85).
@@
$\sup_{k}\,E[\|q_{k}\|]$ $\leq\frac{h_{1}\mu\,L_{1}(n+3)^{3/2}}{2}+h_{1}L_{1}\cdot\frac{(1+2h_{2})h_{1}\sigma}{\sqrt{t}}\sqrt{\frac{2}{1-\rho_{\mathrm{NAG}}}}$
@@
Using [28, Lem. 3], we know that $\|\nabla f(\bar{z}_{k})-\nabla f_{\mu}(\bar{z}_{k})\|\leq\tfrac{\mu}{2}L_{1}(n+3)^{3/2}.$ Thus, the first term in the right-hand side of (88) is bounded and can be manipulated by choosing $h$ and $\mu$ appropriately. For the second term in the right-hand side of (88), we have
@@
where the last inequality uses Jensen’s inequality and steps similar to same as (90). Moreover, we know that $\|v_{k}+l_{k}\|^{2}\leq 2\|v_{k}\|^{2}+2\|l_{k}\|^{2}$ and $E[\|v_{k}\|^{2}]\leq\frac{\sigma^{2}}{t}.$ Similar to the process in (90), we have $E[\|l_{k}\|^{2}]\leq 4L_{1}^{2}E[\|e_{k}\|^{2}]$ and thus
@@
where $\rho=c^{2}+4cL_{1}h+8L_{1}^{2}h^{2}$. For $c\in[0,1)$, to have $\rho<1,$ we require $0<h<\frac{\sqrt{2-c^{2}}-c}{4L_{1}}.$ Since $e_{0}=0,$ by induction we get
$E_{U_{k}}[\|e_{k}\|^{2}]\leq\frac{2h^{2}\sigma^{2}}{t(1-\rho)}=\frac{2h^{2}\sigma^{2}}{t(1-c^{2}-4cL_{1}h-8L_{1}^{2}h^{2})}.$ (95)
@@
$\sup_{k}\|q_{k}\|\leq h\frac{\mu}{2}L_{1}(n+3)^{3/2}+hL_{1}\sqrt{\frac{2h^{2}\sigma^{2}}{t(1-c^{2}-4cL_{1}h-8L_{1}^{2}h^{2})}}$ (96)

### 01 May 2026 v2

@@ 4.2 Binary classification with a neural network
Figure 2 shows the training loss for all six methods (GD, HB, NAG and their ZO counterparts). Shaded regions indicate the standard deviation across 3 independent runs, and the small width confirms a small perturbation. We set $h=1\times 10^{-3}$ for GD and ZO-GD, $h_{1}=1\times 10^{-3}$, $h_{2}=5\times 10^{-4}$ for HB/NAG and their ZO counterparts with smoothing parameter $\mu=10^{-7}$ and number of sampled directions $t=7$. The ZO trajectories closely track their FO counterparts, and all methods reach $99.55\%$ training accuracy and $99.40\%$ test accuracy. It takes approximately 4 times longer to run the ZO methods compared to their FO counterpart. Parameter sensitivity experiments confirming the same phenomena as in the quadratic case are provided in Appendix E.2.
@@
To directly validate the absence of extra dimension dependence, we vary $H\in\{16,32,64,128,256\}$ ($n$ from ${\approx}12{,}700$ to ${\approx}201{,}000$) and run GD and ZO-GD with identical parameters for 1000 iterations. We set $h=5\times 10^{-4},$ $t=10,$ and $\mu=10^{-8}.$ Figure 3  shows that ZO-GD tracks GD at the same rate across all dimensions. Table 1 reports iteration counts to reach within $20\%$ of the FO final loss: the ratio ZO/FO stays close to $1$. Moreover, Table 1 confirms the per-iteration time ratio remains approximately constant.
@@
Table 1: Dimension scaling: iterations to reach within $20\%$ of FO final loss.
@@ 5 Conclusion and future directions
In this work, we revisited the commonly held belief that ZO methods inherently suffer from additional dimension dependence in their iteration complexity compared to their FO counterparts. By adopting a dynamical systems perspective, we showed that, under suitable conditions, the averaged dynamics of ZO algorithms can be interpreted as perturbed versions of their FO counterparts with bounded, controllable perturbations. Leveraging ISS arguments, we proved that ZO methods inherit the same convergence rates as FO methods in expectation, converging to a neighbourhood of the FO fixed point whose size can be made arbitrarily small by tuning the algorithm parameters. We note that while the iteration complexity is dimension-free, achieving a target neighbourhood radius may require parameter choices (e.g., $\mu$ or $t$) that depend on problem parameters, including the variance, so the total function evaluation cost may retain a dimension dependence in the worst case scenario. These findings suggest that the practical efficiency of ZO methods may be closer to that of FO methods than previously understood, particularly when gradient information is unavailable or expensive.
@@ A.1 Comparison with prior work
Table 2 summarises the convergence guarantees of representative ZO methods and highlights how our framework differs from existing analyses. In all prior works, the iteration complexity carries an explicit factor depending on the problem dimension $n$ (or an effective dimension $n_{\mathrm{eff}}$), and convergence is to the exact optimiser. By contrast, our ISS-based framework eliminates the dimension factor from the iteration complexity entirely; the dimension dependence instead appears in the radius of the convergence neighbourhood, which can be made arbitrarily small by tuning the algorithm parameters $\mu$, $h$, and $t$. Moreover, our framework applies uniformly to ZO-GD, ZO-HB, and ZO-NAG, whereas each prior result requires a separate, algorithm-specific analysis. We further note that in prior analyses such as [27], the prescribed step size for ZO-GD scales as $O(1/n)$, which is not consistent with standard FO gradient descent. By contrast, our framework permits the same step size as the FO method, with no dependence on the dimension (see Remark 1).
@@
Reference Assumptions on $f$ Iter. Dim. dependence queries per iter. Convergence target
[27] convex / s.c. $O(n)$ 2 Minimiser
[7] convex $O(n)$ 2 Minimiser
[16] nonconvex $O(n)$ 2 $\|\nabla f\|^{2}\leq\varepsilon$
[39] convex / s.c; Hessian regularity $O(n_{\mathrm{eff}})$ 2 Minimiser
[35] convex; sparse gradients $O(s\log n)$ 2 Minimiser
This work s.c. (Thms. 3–5); n.c + $L_{2}$ reg. (Thm. 6) No explicit $2t$ Neighbourhood of FO fixed point (can be made arbitrarily small)
s.c. = strongly convex; n.c. = non-convex; $2^{\ast}:$ using diminishing step sizes.
@@ Appendix B Additional details and proofs of Section 2
It remains to choose $\bar{r}>0$ and $\bar{q}>0$ so that $\{z_{k}\}_{k\in\mathbb{N}}\subset\mathcal{B}_{R}(z^{e})$. For $z\in\mathbb{R}^{n}$ such that $V(z)\leq\ell=c_{1}R^{2}$, (6) implies that $\|z-z^{e}\|^{2}\leq R^{2}$, i.e., $z\in\mathcal{B}_{R}(z^{e})$, and we can conclude $\{z\in\mathbb{R}^{n}|\ V(z)\leq c_{1}R^{2}\}\subset\mathcal{B}_{R}(z^{e})$. From (41), the set $\{z\in\mathbb{R}^{n}|\ V(z)\leq\ell\}$ is forward-invariant for the perturbed system whenever $\lambda\,\ell+L_{V}\bar{q}\leq\ell$, i.e. whenever $\bar{q}\leq c\,\ell/L_{V}$. Setting
$\bar{r}:=\sqrt{\frac{\ell}{c_{2}}},\qquad\bar{q}:=\frac{c\,\ell}{L_{V}}$
@@
At first glance, $kd$ grows without bound. But recall that $v_{k}$ eventually enters and remains in $0,\sigma(d)+\delta]$ for any fixed $\delta>0$. Let $K^{*}$ be the first time $v_{k}\leq\sigma(d)+\delta$. The strict decrease above $\sigma(d)+\delta$ gives a minimum decrease per step: for $v_{k}\in[\sigma(d)+\delta,\max(R,\sigma(d)+\delta)]$, we have
@@
On the compact set $\sigma(d)+\delta,\max(R,\sigma(d)+\delta)],$ where $\delta>0$ can be arbitrary small (fixed and bounded away from zero), the continuous function $(\mathrm{Id}-\varphi)(s)-d$ is strictly positive (being zero only at $s=\sigma(d)$). Hence, when $R>\sigma(d)+\delta,$ the entry time satisfies the existence of $K^{\ast}$ to have $v_{K^{\ast}}\leq\sigma(d)+\delta$ (and $K^{*}=0$ otherwise). In particular, $K^{*}$ depends only on $R$ and $d$ (not on $v_{0}$ beyond the constraint $v_{0}\leq R$). Now we bound the excess for all $k$:
  • For $k\leq K^{*}$: using $v_{k}\leq\max(R,\sigma(d)+\delta)$ from (42) and $\hat{\beta}(v_{0},k)\geq 0$, we get $e_{k}\leq\max(R,\sigma(d)+\delta)$. But more precisely, by the iterated bound above: $e_{k}=v_{k}-u_{k}\leq K^{*}d$.
  • For $k>K^{*}$: $v_{k}\leq\sigma(d)+\delta$ and $\hat{\beta}(v_{0},k)\geq 0$, so $e_{k}\leq\sigma(d)+\delta$.
@@
$e_{k}\leq\max\bigl(K^{*}d,\;\sigma(d)+\delta\bigr)=:\eta_{R,\delta}(d).$
Since $\delta>0$ is arbitrary, for any $d>0$ we can choose $\delta=d$ (or any class-$\mathcal{K}$ function of $d$), giving
$e_{k}\leq\max\bigl(K^{*}d,\;\sigma(d)+d\bigr)=:\eta_{R}(d).$
Since $\sigma(0)=0$ and $K^{*}=0$ when $d=0$, we have $\eta_{R}(0)=0$. The function $\eta_{R}$ is non-negative, finite, nondecreasing in $d$, and satisfies $\eta_{R}(0)=0$. By regularisation, there exists $\hat{\gamma}_{R}\in\mathcal{K}$ with $\eta_{R}(d)\leq\hat{\gamma}_{R}(d)$ for all $d\geq 0$.
@@ Appendix C Proof of lemmas and theorems of Section 3
Using [27, Lemma 3], we know that $\|\nabla f(\bar{z}_{k})-\nabla f_{\mu}(\bar{z}_{k})\|\leq\tfrac{\mu}{2}L_{1}(n+3)^{3/2}.$ Thus, the first term on the right-hand side (46) is bounded and can be manipulated through the choice of $h$ and $\mu.$ For the second term on the right-hand side of (46), we have
@@
$E[\|e_{k+1}\|]\leq(1+2L_{1}h)E[\|e_{k}\|]+h\frac{\sigma}{\sqrt{t}},$ (49)
@@
$\begin{split}E[\|l_{k}\|^{2}]&=E_{U_{k}}[\|\nabla f_{\mu}(\tilde{z}_{k})-\nabla f_{\mu}(\bar{z}_{k})+\nabla f_{\mu}(\bar{z}_{k})-E_{U_{k}}[\nabla f_{\mu}(\tilde{z}_{k})]\|^{2}]\\ &=E[\|\nabla f_{\mu}(\tilde{z}_{k})-\nabla f_{\mu}(\bar{z}_{k})\|^{2}+\|\nabla f_{\mu}(\bar{z}_{k})-E_{U_{k}}[\nabla f_{\mu}(\tilde{z}_{k})]\|^{2}]\\ &\qquad-2E[\langle\nabla f_{\mu}(\tilde{z}_{k})-\nabla f_{\mu}(\bar{z}_{k}),E_{U_{k}}[\nabla f_{\mu}(\tilde{z}_{k})]-\nabla f_{\mu}(\bar{z}_{k})\rangle]\\ &\leq E[\|\nabla f_{\mu}(\tilde{z}_{k})-\nabla f_{\mu}(\bar{z}_{k})\|^{2}]\\ &\leq L_{1}^{2}E[\|e_{k}\|^{2}].\end{split}$ (53)
@@
For $\beta$-strongly convex $f$ with $L_{1}$-Lipschitz gradients and $h\in\bigl(0,\frac{2}{\beta+L_{1}}\bigr)$, from [28, Thm. 2.1.15] we have
@@
Using [27, Lem. 3], we know that $\|\nabla f_{\mu}(x)-\nabla f(x)\|\leq\frac{\mu}{2}\,L_{1}\,(n+3)^{3/2}$ uniformly in $x$. Thus, the first term in the right-hand side (56) is bounded and can be ensured to be arbitrarily small by selecting $h_{1}$ and $\mu$ appropriately. For the second term in the right-hand side of (56), by $L_{1}$-Lipschitz continuity of $\nabla f$ and Jensen’s inequality, we have
@@
Since we assume that the variance of the random oracle is bounded by $E[\|v_{k}\|^{2}]\leq\sigma^{2}/t$, for $l_{k}$, similar to (53), it holds that $E[\|l_{k}\|^{2}]\leq L_{1}^{2}\,a_{k}$ and thus $E[\|v_{k}+l_{k}\|^{2}]\leq\frac{2\sigma^{2}}{t}+2L_{1}^{2}\,a_{k}$.
@@
$\sup_{k}\,\|q_{k}\|$ $\leq\frac{h_{1}\mu\,L_{1}(n+3)^{3/2}}{2}+h_{1}L_{1}\cdot\frac{h_{1}\sigma}{\sqrt{t}}\,\sqrt{\frac{2}{1-\rho_{\mathrm{HB}}}}$
@@
Using [27, Lem. 3], we know that $\|\nabla f_{\mu}(x)-\nabla f(x)\|\leq\frac{\mu}{2}L_{1}(n+3)^{3/2}$ uniformly in $x$. Thus, the first term in the right-hand side (73) is bounded and can be controlled by the choice of $h_{1}$ and $\mu$. For the second term in the right-hand side of (73), by $L_{1}$-Lipschitz continuity of $\nabla f$ and Jensen’s inequality, we have
@@
Choosing $h_{1}\in(0,\frac{\beta}{2L_{1}^{2}})$ ensures $\psi\in(0,1)$, and then any $h_{2}\in(0,\frac{1}{4}(\psi^{-1/2}-1))$ guarantees (85).
@@
$\sup_{k}\,\|q_{k}\|$ $\leq\frac{h_{1}\mu\,L_{1}(n+3)^{3/2}}{2}+h_{1}L_{1}\cdot\frac{(1+2h_{2})h_{1}\sigma}{\sqrt{t}}\sqrt{\frac{2}{1-\rho_{\mathrm{NAG}}}}$
@@
Using [27, Lem. 3], we know that $\|\nabla f(\bar{z}_{k})-\nabla f_{\mu}(\bar{z}_{k})\|\leq\tfrac{\mu}{2}L_{1}(n+3)^{3/2}.$ Thus, the first term in the right-hand side of (88) is bounded and can be manipulated by choosing $h$ and $\mu$ appropriately. For the second term in the right-hand side of (88), we have
@@
where the last inequality uses Jensen’s inequality and steps similar to same as (90). Moreover, we know that $\|v_{k}+l_{k}\|^{2}\leq 2\|v_{k}\|^{2}+2\|l_{k}\|^{2}$ and $E[\|v_{k}\|^{2}]\leq\frac{\sigma^{2}}{t}.$ Similar to the process in (90) and considering the fact that $E[\|X-E[X]\|^{2}]\leq E[\|X\|^{2}]$, we have $E[\|l_{k}\|^{2}]\leq L_{1}^{2}E[\|e_{k}\|^{2}]$ and thus
@@
where $\rho=c^{2}+4cL_{1}h+2L_{1}^{2}h^{2}$. For $c\in[0,1)$, to have $\rho<1,$ we require $h\in\bigl(0,\frac{\sqrt{2(1+c^{2})}-2c}{2L_{1}}\bigr).$ Since $e_{0}=0,$ by induction we get
$E_{U_{k}}[\|e_{k}\|^{2}]\leq\frac{2h^{2}\sigma^{2}}{t(1-\rho)}=\frac{2h^{2}\sigma^{2}}{t(1-c^{2}-4cL_{1}h-2L_{1}^{2}h^{2})}.$ (95)
@@
$\sup_{k}\|q_{k}\|\leq h\frac{\mu}{2}L_{1}(n+3)^{3/2}+hL_{1}\sqrt{\frac{2h^{2}\sigma^{2}}{t(1-c^{2}-4cL_{1}h-2L_{1}^{2}h^{2})}}$ (96)